# Implicit-Explicit Trust Region Method for Computing Second-Order Stationary Points of a class of Landau Models

### 04 Mar 2026 v1

@@
Abstract.
We propose an implicit-explicit trust region method for computing second-order stationary points of a class of Landau-type free energy functionals, which correspond to physically (meta-)stable phases. The proposed method is demonstrated through the Landau-Brazovskii (LB) model in this work, while broader applicability to more Landau models of the similar type is straightforwardly extended. The LB energy functional is discretized via the Fourier pseudospectral method, which yields a finite-dimensional nonconvex optimization problem. By exploiting the Hessian structure—–specifically, that the interaction potential is diagonal in reciprocal space whereas the bulk energy is diagonal in physical space—–we design an adaptive implicit-explicit solver for the trust region subproblem. This solver utilizes the fast Fourier transform to perform efficient matrix-vector products, significantly reducing computational complexity while ensuring provable convergence to the global minimizer of the subproblem. In contrast to existing algorithms that target first-order stationary points, our proposed method can converge to a second-order stationary state, corresponding to a local minimum with theoretical convergence guarantees. Numerical experiments on the LB model demonstrate that the proposed approach efficiently escapes saddle points and significantly outperforms existing first-order schemes. Furthermore, we successfully identify the stable region of the FDDD phase, a structure previously unreported in the LB phase diagram.
2020 Mathematics Subject Classification:
Primary 65N22, 65K10, 90C26
The first author was supported by the National Natural Science Foundation of China (12271291) and New Cornerstone Investigator Program (100001127). The third author was supported by the National Key R&D Program of China (2023YFA1008802), the National Natural Science Foundation of China (12288101), the Science and Technology Innovation Program of Hunan Province (2024RC1052). The fourth author was supported by the National Key R&D Program of China (2023YFB3001604). This work is also partially supported by the High Performance Computing Platform of Xiangtan University.
1. Introduction
Landau theory constitutes a foundational framework for describing phases and phase transitions across a wide range of physical systems [22, 21]. In this paper, we consider a class of Landau-type energy functionals that contain two components, an interaction potential and a bulk energy term,
@@ 1. Introduction
where $\psi(\bm{x})$ is the order parameter. In this formulation, the interaction potential $G(\psi)$ typically involves differential or non-local operators that describe spatial correlations and pattern selection, while the bulk energy term $F(\psi)$ is a nonlinear potential driving phase transitions. This unified framework encompasses a wide variety of physical models, such as the Ginzburg-Landau model, the Ohta-Kawasaki model, the Swift-Hohenberg model, the LB model, and multi-length-scale models [17, 25, 39, 31, 35]. One of the most important problems in investigating these models is computing both stable and metastable states accurately and efficiently [13, 24], which is crucial for constructing phase diagrams and guiding the design of ordered materials in modulated systems.
@@
SP-Is encompass global minima, local minima, and saddle points (SDPs), corresponding to stable, metastable and transition states, respectively. SP-IIs form a subset of SP-Is and represent physically stable and metastable states. The identification of SP-IIs enables accurate determination of equilibrium states. We define SP-IIs using positive semi-definite Hessians, rather than strictly positive definite ones, to accommodate degenerate stationary states with zero eigenvalues in their Hessian matrices. Such degeneracy arises fundamentally from continuous symmetries, most notably translational invariance in ordered phases including periodic crystals and quasicrystals [3, 13, 14].
The current algorithms that aim to compute stationary states for such Landau models can be broadly classified into three categories. The first category is to solve the steady nonlinear Euler–Lagrange equations of (1.1). The second category focuses on solving the nonlinear gradient flow equation. Approaches include explicit-implicit schemes [36, 11, 33], the scalar auxiliary variable [32], the invariant energy quadrature scheme [38], and exponential time differencing methods [16]. The third category is from the optimization viewpoint. These approaches directly find the stationary states via solving a finite-dimensional or infinite-dimensional nonconvex optimization problem, such as gradient descent [28, 34], proximal gradient, and the Bregman proximal gradient approaches [3, 4, 18]. However, these approaches typically only consider convergence to SP-Is and theoretically cannot guarantee convergence to SP-IIs.
To address this limitation, this work develops an algorithm that can provably converge to SP-IIs with theoretical guarantees by combining recent advances in optimization. In particular, recent theories suggest that SP-IIs in nonconvex problems can be reached either through controlled perturbations or by exploiting curvature information. Stochastic gradient descent and its adaptive forms escape from saddle points and converge to SP-IIs by introducing noise during the iteration [20, 23]. Trust region (TR) methods [15, 29], and the variants of Newton’s method [27, 30, 1] can also ensure convergence to SP-IIs by utilizing the Hessian information.
To validate the effectiveness of the proposed algorithm, we employ the LB model as a primary benchmark. This model is a representative example that describes systems exhibiting spatially modulated phases at finite wavenumbers, such as block copolymers, liquid crystals, microemulsions, and magnetic materials [7, 35, 8]. It effectively captures a broad spectrum of ordered phases, including lamellar, hexagonal, body-centered cubic spheres, face-centered cubic spheres, double gyroid, A15, and $\sigma$ phase [26]. The energy functional of the LB model is
@@
Notation: In the rest of the content, let $\|\bm{u}\|_{2}=\sqrt{\bm{u}^{\top}\bm{u}}$ denote the $\ell_{2}$ norm for vector $\bm{u}$. The operator (spectral) norm of a matrix $\bm{H}$ is denoted by $\|\bm{H}\|:=\max\limits_{\|\bm{u}\|_{2}=1}\|\bm{H}\bm{u}\|_{2}$. For a third-order tensor $\mathcal{T}$ (e.g., third-order derivative $\nabla^{3}E$), the operator norm is defined as $\|\mathcal{T}\|:=\max\limits_{\|\bm{u}\|_{2}=\|\bm{v}\|_{2}=\|\bm{w}\|_{2}=1}|\mathcal{T}[\bm{u},\bm{v},\bm{w}]|$, which represents the maximum absolute value of the tensor evaluated over all unit vectors [10].
2. Discretizing the LB model with Fourier pseudospectral method
@@
3. Implicit-Explicit Trust Region Method
@@
3.1. IMEX-TR Method and Convergence Analysis
@@
Theorem 3.1 ([12], Corollary 7.2.2).
@@
3.1.1. Convergence to Second-Order Stationary Points
@@
where $h_{\text{Lip}}:=\sup_{\bm{w}\in\bar{\mathcal{B}}}\|\nabla^{3}\bar{E}(\bm{w})\|<\infty$. This completes the proof. ∎
@@
Proof.
In combination with Theorem 3.2, the properties of the sequence $\{\bm{v}_{j}\}$ follow directly from Theorem 3.14 and Theorem 3.26 of [15]. ∎
3.2. Efficient Global Solver for the Nonconvex TR Subproblem
@@
3.2.1. Global Convergence of the IMEX solver for the TR subproblem
@@
The second equality above is derived from Corollary 3.5. ∎
@@
The second equality above follows from Corollary 3.5 and $\beta_{k+1}^{2}=\lambda_{k+1}$, (3.5) has been used for the fourth equality. Let $b=1/\eta_{\min}+\|\bm{T}\|$ finishes the proof. ∎
Using the KŁ property and the techniques of Lemma 2.6 from [2], we prove that the sequence $\{\bm{d}_{k}\}$ is a Cauchy sequence.
Definition 3.8 (KŁ Property[5]).
@@
Proof.
As $\{\bm{z}_{k}\}$ is bounded, there must exist an accumulation point of $\{\bm{z}_{k}\}$, let it be $\bm{z}_{*}=(\bm{d}_{*},\beta_{*})$, and there exists a subsequence $\{\bm{z}_{k_{l}}\}$ of $\{\bm{z}_{k}\}$ converging to $\bm{z}_{*}$ as $l\to\infty$. Due to the boundedness from below of $L$ and the monotonicity of $\{L(\bm{z}_{k})\}$, the continuous property of $L$ implies that $\{L(\bm{z}_{k})\}$ converges to $L(\bm{z}_{*})$ as $k\to\infty$ with $L(\bm{z}_{*})<L(\bm{z}_{k})$, $k\in\mathbb{N}$. In particular, $L$ satisfies the KŁ property [5], and thus
@@
which implies that the sequence $\{\bm{d}_{k}\}$ is a Cauchy sequence, which therefore converges to $\bm{d}_{*}$ since it is an accumulation point of $\{\bm{d}_{k}\}$. ∎
@@
combining with Corollary 3.5, we conclude that $(\bm{d}_{*},\lambda_{*})$ is a KKT point of problem (3.4). ∎
@@
Proposition 3.11 (Proposition 2.2 [9]).
@@
it is guaranteed that $d^{(1)}_{k+1}g^{(1)}\leq 0$. Thus, by induction, (3.18) holds. Let $k\to\infty$ in (3.18) and using Proposition 3.11, we conclude that $\bm{d}_{*}$ must be a global minimizer of problem (3.4). ∎
4. Numerical Experiments
@@
4.1. Convergence
In this subsection, we compare the IMEX-TR method with representative methods, including the adaptive accelerated Bregman proximal gradient (AA-BPG) [18], and several gradient flow methods, including the semi-implicit scheme (SIS), first-order stabilized semi-implicit scheme (SSIS1), second-order stabilized semi-implicit scheme (SSIS2) [33], the invariant energy quadratization (IEQ) [38], the scalar auxiliary variable (SAV) [32], and exponential time differencing (ETD) methods [16]. [^1] Since the results of SIS, SSIS1, and IEQ are similar to those of other methods except IMEX-TR, the corresponding results are not reported. All algorithms start from the same initial condition to demonstrate that the IMEX-TR method, by effectively using Hessian information, converges to SP-IIs, while other methods may converge to SP-Is.
We illustrate the performance of IMEX-TR through three different examples. Table 1 summarizes the model parameters $(\tau,\gamma)$, computational domains, and initial configurations employed in the three numerical experiments. The initial ordered configurations—specifically, three-dimensional lamellar (LAM) and hexagonal (HEX) phases—are constructed via Fourier series expansion of the form [19]
@@ 4.1. Convergence
Table 1. Model parameters, computational domains, and initial configurations for three numerical examples.
@@
Table 2. Initial reciprocal lattice point sets $\Xi$, nonzero Fourier coefficients $\hat{\psi}(\bm{k})$, and reciprocal lattice matrices $\bm{B}$ for the LAM and HEX phases.
@@
Figure 1 presents the converged SPs and their corresponding energy values. In Examples 1 and 2, conventional first-order schemes fail to escape the initial energy basins and stagnate at unstable stationary points. Specifically, in Example 1, first-order methods remain trapped in the LAM state, while in Example 2, they converge to a disordered state (i.e., $\phi(\bm{x})\approx 0$). In contrast, the IMEX-TR method successfully converges to SP-IIs, HEX and body-centered cubic (BCC) phases, both exhibiting lower free energies. In Example 3, all methods reach the SP-II state HEX; however, the IMEX-TR method escapes the HEX basin and discovers a cubic FDDD pattern that achieves an energy of $E=-8.05\times 10^{-2}$, which is markedly lower than the HEX phase ($E=-8.02\times 10^{-2}$). The cubic FDDD pattern has been reported in the phase diagram of AB-type diblock copolymers [37]; however, its presence in the LB model has not been previously documented.
@@
Figure 1. Final physical phases and energy. IMEX-TR identifies the (meta-)stable HEX, BCC, and cubic FDDD phases across the three examples, consistently achieving lower energy states compared to standard first-order schemes.
Figure 2. Evolution of the four smallest eigenvalues of the Hessian matrix. The dashed line indicates the zero threshold. IMEX-TR consistently reaches the SP-II region (non-negative eigenvalues), whereas first-order methods often stagnate at SDPs with negative curvature.
4.2. Robustness to initial conditions
@@
Figure 3. Trajectories of IMEX-TR activated at different stages of first-order methods. Whether started from the initial point, an intermediate state, or the converged saddle point, IMEX-TR consistently converges to the same (meta-)stable SP-II (HEX phase), demonstrating its efficacy in exploiting negative curvature directions to escape unstable stationary points.
4.3. Construction of new phase diagram
The phase diagram of the LB model has been well documented to contain stable ordered phases such as LAM, HEX, BCC, double gyroid (DG), face-centered cubic spheres, A15, and $\sigma$ phases [6, 26]. Our simulation in Section 4.1 demonstrates that the cubic FDDD pattern is indeed an SP-II according to the IMEX-TR method. There is a stable region of the discovered cubic FDDD pattern in the phase diagram of AB-type diblock copolymers [37]. However, its thermodynamic stability in the LB model remains uncertain. Motivated by this finding, we conducted an extensive search across the $(\tau,\gamma)$ parameter space to investigate whether a stable region for the cubic FDDD pattern exists within the LB model.
Figure 4. (a) Phase diagram of the LB model. The purple shaded area corresponds to the stable region of the FDDD phase. (b)-(f) illustrate the SP-II phases: LAM, HEX, BCC, DG, and FDDD.
As presented in Figure 4, our results confirm the existence of a stable region for the cubic FDDD pattern. Moreover, the identified stable region of the cubic FDDD phase does not affect the stable regions of the face-centered cubic spheres, A15, and $\sigma$ phases [26]. This finding not only provides a more complete picture of the LB phase diagram but also demonstrates the capability of IMEX-TR to explore complex energy landscapes and uncover new stable structures.
5. Conclusion

### 14 Mar 2026 v2

@@
Abstract
We propose an implicit-explicit trust region method for computing second-order stationary points of a class of Landau-type free energy functionals, which correspond to physically (meta-)stable phases. The proposed method is demonstrated through the Landau-Brazovskii (LB) model in this work, while broader applicability to more Landau models of the similar type is straightforwardly extended. The LB energy functional is discretized via the Fourier pseudospectral method, which yields a finite-dimensional nonconvex optimization problem. By exploiting the Hessian structure—specifically, that the interaction potential is diagonal in reciprocal space whereas the bulk energy is diagonal in physical space—we design an adaptive implicit-explicit solver for the trust region subproblem. This solver utilizes the fast Fourier transform to perform efficient matrix-vector products, significantly reducing computational complexity while ensuring provable convergence to the global minimizer of the subproblem. In contrast to existing algorithms that target first-order stationary points, our proposed method can converge to a second-order stationary state, corresponding to a local minimum with theoretical convergence guarantees. Numerical experiments on the LB model demonstrate that the proposed approach efficiently escapes saddle points and significantly outperforms existing first-order schemes. Furthermore, we successfully identify the stable region of the FDDD phase, a structure previously unreported in the LB phase diagram.
2020 MSC. 65N22, 65K10, 90C26.
1 Introduction
Landau theory constitutes a foundational framework for describing phases and phase transitions across a wide range of physical systems [landau1936theory, landau1958statistical]. In this paper, we consider a class of Landau-type energy functionals that contain two components, an interaction potential and a bulk energy term,
@@ 1 Introduction
where $\psi(\bm{x})$ is the order parameter. In this formulation, the interaction potential $G(\psi)$ typically involves differential or non-local operators that describe spatial correlations and pattern selection, while the bulk energy term $F(\psi)$ is a nonlinear potential driving phase transitions. This unified framework encompasses a wide variety of physical models, such as the Ginzburg-Landau model, the Ohta-Kawasaki model, the Swift-Hohenberg model, the LB model, and multi-length-scale models [ginzburg2009theory, liu2003phase, zhang2008efficient, savitz2018multiple, swift1977hydrodynamic]. One of the most important problems in investigating these models is computing both stable and metastable states accurately and efficiently [cui2025efficient, lin2010numerical], which is crucial for constructing phase diagrams and guiding the design of ordered materials in modulated systems.
@@
SP-Is encompass global minima, local minima, and saddle points (SDPs), corresponding to stable, metastable and transition states, respectively. SP-IIs form a subset of SP-Is and represent physically stable and metastable states. The identification of SP-IIs enables accurate determination of equilibrium states. We define SP-IIs using positive semi-definite Hessians, rather than strictly positive definite ones, to accommodate degenerate stationary states with zero eigenvalues in their Hessian matrices. Such degeneracy arises fundamentally from continuous symmetries, most notably translational invariance in ordered phases including periodic crystals and quasicrystals [bao2024convergence, cui2025efficient, cui2024spring].
The current algorithms that aim to compute stationary states for such Landau models can be broadly classified into three categories. The first category is to solve the steady nonlinear Euler–Lagrange equations of (1.1). The second category focuses on solving the nonlinear gradient flow equation. Approaches include explicit-implicit schemes [wise2009energy, chen1998applications, shen2010numerical], the scalar auxiliary variable [shen2019new], the invariant energy quadrature scheme [yang2016linear], and exponential time differencing methods [fu2022energy]. The third category is from the optimization viewpoint. These approaches directly find the stationary states via solving a finite-dimensional or infinite-dimensional nonconvex optimization problem, such as gradient descent [neuberger2009sobolev, sial2003energy], proximal gradient, and the Bregman proximal gradient approaches [bao2024convergence, bao2020adaptive, jiang2020efficient]. However, these approaches typically only consider convergence to SP-Is and theoretically cannot guarantee convergence to SP-IIs.
To address this limitation, this work develops an algorithm that can provably converge to SP-IIs with theoretical guarantees by combining recent advances in optimization. In particular, recent theories suggest that SP-IIs in nonconvex problems can be reached either through controlled perturbations or by exploiting curvature information. Stochastic gradient descent and its adaptive forms escape from saddle points and converge to SP-IIs by introducing noise during the iteration [jin2017escape, lee2016gradient]. Trust region (TR) methods [curtis2017trust, nouiehed2020trust], and the variants of Newton’s method [nesterov2006cubic, royer2020newton, agarwal2017finding] can also ensure convergence to SP-IIs by utilizing the Hessian information.
To validate the effectiveness of the proposed algorithm, we employ the LB model as a primary benchmark. This model is a representative example that describes systems exhibiting spatially modulated phases at finite wavenumbers, such as block copolymers, liquid crystals, microemulsions, and magnetic materials [brazovskiui1996phase, swift1977hydrodynamic, caplan2017]. It effectively captures a broad spectrum of ordered phases, including lamellar, hexagonal, body-centered cubic spheres, face-centered cubic spheres, double gyroid, A15, and $\sigma$ phase [mcclenagan2019landau]. The energy functional of the LB model is
@@
Notation: In the rest of the content, let $\|\bm{u}\|_{2}=\sqrt{\bm{u}^{\top}\bm{u}}$ denote the $\ell_{2}$ norm for vector $\bm{u}$. The operator (spectral) norm of a matrix $\bm{H}$ is denoted by $\|\bm{H}\|:=\max\limits_{\|\bm{u}\|_{2}=1}\|\bm{H}\bm{u}\|_{2}$. For a third-order tensor $\mathcal{T}$ (e.g., third-order derivative $\nabla^{3}E$), the operator norm is defined as $\|\mathcal{T}\|:=\max\limits_{\|\bm{u}\|_{2}=\|\bm{v}\|_{2}=\|\bm{w}\|_{2}=1}|\mathcal{T}[\bm{u},\bm{v},\bm{w}]|$, which represents the maximum absolute value of the tensor evaluated over all unit vectors [cartis2020sharp].
2 Discretizing the LB model with Fourier pseudospectral method
@@
3 Implicit-Explicit Trust Region Method
@@
3.1 IMEX-TR Method and Convergence Analysis
@@
Theorem 3.1 ([conn2000trust], Corollary 7.2.2).
@@
3.1.1 Convergence to Second-Order Stationary Points
@@
where $h_{\text{Lip}}:=\sup_{\bm{w}\in\bar{\mathcal{B}}}\|\nabla^{3}\bar{E}(\bm{w})\|<\infty$. This completes the proof.
@@
Proof 3.5.
In combination with Theorem 3.2, the properties of the sequence $\{\bm{v}_{j}\}$ follow directly from Theorem 3.14 and Theorem 3.26 of [curtis2017trust].
3.2 Efficient Global Solver for the Nonconvex TR Subproblem
@@
3.2.1 Global Convergence of the IMEX solver for the TR subproblem
@@
The second equality above is derived from Corollary 3.7.
@@
The second equality above follows from Corollary 3.7 and $\beta_{k+1}^{2}=\lambda_{k+1}$, (3.5) has been used for the fourth equality. Let $b=1/\eta_{\min}+\|\bm{T}\|$ finishes the proof.
Using the KŁ property and the techniques of Lemma 2.6 from [attouch2013convergence], we prove that the sequence $\{\bm{d}_{k}\}$ is a Cauchy sequence.
Definition 3.12 (KŁ Property[bolte2014proximal]).
@@
Proof 3.14.
As $\{\bm{z}_{k}\}$ is bounded, there must exist an accumulation point of $\{\bm{z}_{k}\}$, let it be $\bm{z}_{*}=(\bm{d}_{*},\beta_{*})$, and there exists a subsequence $\{\bm{z}_{k_{l}}\}$ of $\{\bm{z}_{k}\}$ converging to $\bm{z}_{*}$ as $l\to\infty$. Due to the boundedness from below of $L$ and the monotonicity of $\{L(\bm{z}_{k})\}$, the continuous property of $L$ implies that $\{L(\bm{z}_{k})\}$ converges to $L(\bm{z}_{*})$ as $k\to\infty$ with $L(\bm{z}_{*})<L(\bm{z}_{k})$, $k\in\mathbb{N}$. In particular, $L$ satisfies the KŁ property [bolte2014proximal], and thus
@@
which implies that the sequence $\{\bm{d}_{k}\}$ is a Cauchy sequence, which therefore converges to $\bm{d}_{*}$ since it is an accumulation point of $\{\bm{d}_{k}\}$.
@@
combining with Corollary 3.7, we conclude that $(\bm{d}_{*},\lambda_{*})$ is a KKT point of problem (3.4).
@@
Proposition 3.17 (Proposition 2.2 [carmon2020first]).
@@
it is guaranteed that $d^{(1)}_{k+1}g^{(1)}\leq 0$. Thus, by induction, (3.18) holds. Let $k\to\infty$ in (3.18) and using Proposition 3.17, we conclude that $\bm{d}_{*}$ must be a global minimizer of problem (3.4).
4 Numerical Experiments
@@
4.1 Convergence
In this subsection, we compare the IMEX-TR method with representative methods, including the adaptive accelerated Bregman proximal gradient (AA-BPG) [jiang2020efficient], and several gradient flow methods, including the semi-implicit scheme (SIS), first-order stabilized semi-implicit scheme (SSIS1), second-order stabilized semi-implicit scheme (SSIS2) [shen2010numerical], the invariant energy quadratization (IEQ) [yang2016linear], the scalar auxiliary variable (SAV) [shen2019new], and exponential time differencing (ETD) methods [fu2022energy]. [^1] Since the results of SIS, SSIS1, and IEQ are similar to those of other methods except IMEX-TR, the corresponding results are not reported. All algorithms start from the same initial condition to demonstrate that the IMEX-TR method, by effectively using Hessian information, converges to SP-IIs, while other methods may converge to SP-Is.
We illustrate the performance of IMEX-TR through three different examples. Table 1 summarizes the model parameters $(\tau,\gamma)$, computational domains, and initial configurations employed in the three numerical experiments. The initial ordered configurations—specifically, three-dimensional lamellar (LAM) and hexagonal (HEX) phases—are constructed via Fourier series expansion of the form [jiang2013discovery]
@@ 4.1 Convergence
Table 1: Model parameters, computational domains, and initial configurations for three numerical examples.
@@
Table 2: Initial reciprocal lattice point sets $\Xi$, nonzero Fourier coefficients $\hat{\psi}(\bm{k})$, and reciprocal lattice matrices $\bm{B}$ for the LAM and HEX phases.
@@
Figure 1 presents the converged SPs and their corresponding energy values. In Examples 1 and 2, conventional first-order schemes fail to escape the initial energy basins and stagnate at unstable stationary points. Specifically, in Example 1, first-order methods remain trapped in the LAM state, while in Example 2, they converge to a disordered state (i.e., $\phi(\bm{x})\approx 0$). In contrast, the IMEX-TR method successfully converges to SP-IIs, HEX and body-centered cubic (BCC) phases, both exhibiting lower free energies. In Example 3, all methods reach the SP-II state HEX; however, the IMEX-TR method escapes the HEX basin and discovers a cubic FDDD pattern that achieves an energy of $E=-8.05\times 10^{-2}$, which is markedly lower than the HEX phase ($E=-8.02\times 10^{-2}$). The cubic FDDD pattern has been reported in the phase diagram of AB-type diblock copolymers [yamada2006fddd]; however, its presence in the LB model has not been previously documented.
@@
Figure 1: Final physical phases and energy. IMEX-TR identifies the (meta-)stable HEX, BCC, and cubic FDDD phases across the three examples, consistently achieving lower energy states compared to standard first-order schemes.
Figure 2: Evolution of the four smallest eigenvalues of the Hessian matrix. The dashed line indicates the zero threshold. IMEX-TR consistently reaches the SP-II region (non-negative eigenvalues), whereas first-order methods often stagnate at SDPs with negative curvature.
4.2 Robustness to initial conditions
@@
Figure 3: Trajectories of IMEX-TR activated at different stages of first-order methods. Whether started from the initial point, an intermediate state, or the converged saddle point, IMEX-TR consistently converges to the same (meta-)stable SP-II (HEX phase), demonstrating its efficacy in exploiting negative curvature directions to escape unstable stationary points.
4.3 Construction of new phase diagram
The phase diagram of the LB model has been well documented to contain stable ordered phases such as LAM, HEX, BCC, double gyroid (DG), face-centered cubic spheres, A15, and $\sigma$ phases [brazovskii1975phase, mcclenagan2019landau]. Our simulation in Section 4.1 demonstrates that the cubic FDDD pattern is indeed an SP-II according to the IMEX-TR method. There is a stable region of the discovered cubic FDDD pattern in the phase diagram of AB-type diblock copolymers [yamada2006fddd]. However, its thermodynamic stability in the LB model remains uncertain. Motivated by this finding, we conducted an extensive search across the $(\tau,\gamma)$ parameter space to investigate whether a stable region for the cubic FDDD pattern exists within the LB model.
Figure 4: (a) Phase diagram of the LB model. The purple shaded area corresponds to the stable region of the FDDD phase. (b)-(f) illustrate the SP-II phases: LAM, HEX, BCC, DG, and FDDD.
As presented in Figure 4, our results confirm the existence of a stable region for the cubic FDDD pattern. Moreover, the identified stable region of the cubic FDDD phase does not affect the stable regions of the face-centered cubic spheres, A15, and $\sigma$ phases [mcclenagan2019landau]. This finding not only provides a more complete picture of the LB phase diagram but also demonstrates the capability of IMEX-TR to explore complex energy landscapes and uncover new stable structures.
5 Conclusion